Find the minimum amount of tin sheet in square inches that can be made into a closed cylinder having a volume of 108 cubic inches


1
Expert's answer
2021-06-14T15:38:54-0400

"V=\pi R^2L=108"


"L=108/\pi R^2"


"S=2\pi RL+2\pi R^2=216/R+2\pi R^2"


"\frac{dS}{dR}=-216/R^2+4\pi R=0"


"4\pi R^3-216=0"


"R^3=216/(4\pi)"


"R_{min}=2.58\ in"


"S_{min}=216/2.58+2\pi\cdot2.58^2=125.53\ in^2"